# GPU Acceleration of SQL Analytics on Compressed Data

### v1

@@ 9.2. Production queries
Figure 10 shows the sizes for the in-memory representation of the 15 columns of the fact table. The series ‘Plain’ refers to the default Plain encoding (with dictionary encoding) but with no other compression schemes applied. The total size for the 15 columns is 120.36 GiB with Plain and does not fit into the 80 GiB GPU HBM. Considering only the subset of columns needed by the specific queries, Q1 needs 87.53 GiB and Q2 and Q3 need 82.06 GiB, which is still larger than the GPU HBM capacity. The bottom row of the x-axis shows the encoding scheme for Compressed—‘C’: Composite (Plain + Index), ‘R’: RLE, ‘P’: Plain with bit-width reduction based on the range of values. The total size for 15 columns with Compressed is 56.84 GiB. For Q1 columns it is 43.69 GiB and for Q2 and Q3 it is 30.86 GiB, which can fit in the GPU HBM and also leave space for intermediate results. For reference, the on-disk V-ordered file size for the projected fact table for the 15 columns is 29.87 GiB.
This production dataset has substantial opportunities for RLE compression. Figure 11 shows the number of runs and average run lengths in 7 of the 15 fact table columns that were compressed using RLE. Column 7 has a single run of 2.94B rows (same value for all rows). Average run lengths for other columns range from several tens to hundreds of thousands. For example, column 11 has 85.35M runs with an average run length of 34.41. Each run needs 20 bytes for representation—4 for value ($v$), 8 each for start ($s$) and end ($e$) positions. This requires 1.59 GiB memory for the column, which is significantly less than 10.94 GiB (= 4 x 2.94B) needed for the Plain representation. Further savings are possible for RLE by using narrower data types for $s$ and $e$.
Figure 12 shows query run times on state-of-the-art CPU-based commercial RDBMS, Microsoft SQL Server, and analytical data engine, Analysis Services. Our techniques lead to substantial speedups—15.75 (Q1), 13.66 (Q2), 9.82 (Q3), 12.76 (total) over SQL Server and 8.72 (Q1), 13.16 (Q2), 8.05 (Q3), 9.52 (total) over Analysis Services.
Finally, we study scalability characteristics by varying the input data size. For this we create smaller datasets with the first 5% (0.15B), 20% (0.6B), and 50% (1.5B) rows of the fact table. Figure 13 shows good performance scalability since GPU query times on compressed data increase much slower with data size compared to GPU query times on plain data. Queries cannot be run on Plain data for the 50% (and 100%) dataset due to running out of GPU HBM capacity, and Q2 and Q3 GPU times for just the 20% dataset exceed that of SQL Server and Analysis Services on CPUs for the 100% dataset! Thus, query execution on compressed data is necessary to unlock the GPU acceleration potential for this workload.
We could fit and process even larger datasets in the GPU HBM with our compression schemes. Figure 14 shows actuals and projections, using a linear model, of peak GPU memory consumption during query runs as we increase the fact table size. For Plain, we can only fit less than 50% of fact table rows whereas with Compressed we could go up to 157% (4.62B) for Q1, 222% (6.52B) for Q2, and 216% (6.36B) for Q3.
10. Related Work
GPU acceleration for SQL analytics: There has been considerable research into using GPUs to accelerate SQL analytical queries (Shanbhag et al., 2020; Cao et al., 2023; He et al., 2022; Yuan et al., 2013; Rosenfeld et al., 2022; Yogatama et al., 2022; Sioulas et al., 2019; Yogatama et al., 2025; Yuan et al., 2025; Chrysogelos et al., 2019; Wu et al., 2025; Hu et al., 2022; Rui and Tu, 2017; Hong et al., 2025; Deng et al., 2024; Karnagel et al., 2015; HeavyDB, ssed; Data, 2025; Paul et al., 2021; BlazingSQL, 2021; Lutz et al., 2020; Maltenberger et al., 2022; Rui et al., 2020). Rapid advances in GPU technology, to meet the needs of AI applications, continue to make GPUs increasingly attractive for SQL acceleration. Our work adds further support for this paradigm by developing new methods for GPU query processing on light-weight compressed data.
11. Conclusion

### v2

@@ C.3. Scalability Analysis
Figure 19 shows actuals and projections, using a linear model, of peak GPU memory consumption during query runs as we increase the fact table size. For Plain, we can only fit less than 50% of fact table rows whereas with Compressed we could go up to 157% (4.62B) for Q1, 222% (6.52B) for Q2, and 216% (6.36B) for Q3.
@@
Appendix D Query Optimization Rules
Our query execution plans are based on optimized versions of the tensor program query plans from TQP, with the following optimizations applied:
  • Apply predicates to RLE columns before Plain columns: RLE columns can be filtered more efficiently due to their compressed representation, and they are potentially highly selective. When an RLE predicate is selective, it produces an Index mask that can be applied to Plain columns, avoiding the need to evaluate the Plain predicate on all rows.
  • Composite predicate evaluation on RLE columns: When an RLE column has multiple predicates (e.g., col > 10 AND col < 100), we evaluate all predicates directly on the RLE value tensor to produce a single boolean mask, then apply it to the start and end positions once. This avoids the overhead of creating intermediate RLE masks and repeatedly applying the range intersection algorithm (Algorithm 1). This is in contrast with Plain columns, where predicates are evaluated one-by-one to produce individual masks that are then combined with AND operations, as Plain mask AND operations are much less costly than RLE range intersections.
  • Join ordering to prioritize RLE join columns: If a table participates in multiple join and/or semi-join operations, we do the joins/semi-joins involving the RLE columns before the Plain columns. This helps to avoid/reduce fragmentation of RLE runs that could otherwise happen if the joins/semi-joins are performed with Plain columns first. In the spirit of predicate pushdown, prioritizing RLE joins allows early filtering of entire runs when join conditions are not satisfied, reducing the data volume for subsequent operations.
  • Avoiding redundant filter operations in groupby-aggregate: If in a filter-groupby-aggregate query pattern, the groupby columns are RLE, then we apply the filter operations just on the groupby columns and not on the aggregate columns. This is because the RLE groupby columns will track the filtered ranges through the adjusted start and end positions, and will automatically only consider rows for aggregation within those ranges. Thus, explicitly applying the filter operations on the aggregate columns as well is functionally correct, but redundant.
We find that these optimizations also benefit Plain processing, explaining why our Plain implementation outperforms baseline TQP: (1) selection pushdown - evaluating selections and (semi-)joins over potentially highly selective columns first produces masks that reduce work on other columns, and (2) NULL optimization - our query plans exclude NULL handling operators and assume non-NULL data, allowing count(*) to be computed once for all aggregations rather than per-column counts. We acknowledge that the lack of NULL support is a limitation. These optimizations are currently applied manually; incorporating them into a data encoding-aware query optimizer and supporting NULL handling in compressed representations are left as future work.